OpenStudy (anonymous):

What is the total cost to repay a $150,000 mortgage with a 5% interest rate for a 30-year term? Please help will give a medal

OpenStudy (tkhunny):

Do you have a formula for the periodic repayment of a loan over a specified period? Mortgages are usually monthly.

OpenStudy (tkhunny):

i = 0.05 j = i/12 = 0.00416666... v = 1/(1-j) = 0.9958506225... -- Monthly n = 20 years m = 12 payments annually T = Term = n*m = 360 months \(150000 = Pmt\cdot\dfrac{1−v^{T}}{j}\) Since we know v and T and j, we can solve for Pmt. Then, we can do anything the question asks of us. You should know this formula inside and out. Better yet, you should be able to derive it at the drop of a hat. You WILL be required to know it on an exam.

OpenStudy (anonymous):

thanks

OpenStudy (tkhunny):

It is time for you to start showing YOUR work.